OpenStudy (anonymous):

Anyone wanna help me? I'm having trouble finding the first payment :/ Jill Barkely obtained a 25-year, $460,000 mortgage loan from University Savings and Loan Association with 6% interest. The monthly payment is $2,962.40. For the first payment, find the interest to the nearest cent.

OpenStudy (tkhunny):

Have you considered calculating it directly? $2,962.40*(0.06/12)

OpenStudy (anonymous):

which is 14.81 correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I tried that but it was wrong

OpenStudy (tkhunny):

Did you try $14.82? There are some rounding issues. For one, $2,962.40 gets only to $459,784.81, a little short of the suggested amount.

OpenStudy (anonymous):

Yes it is still incorrect :/ I also tried this too but I'm not sure if I calculated correctly. Originally I had gotten $7.60

OpenStudy (tkhunny):

It doesn't matter how you calculate it. The answer is unique.

OpenStudy (tkhunny):

\(2962.40\cdot\dfrac{1-\left(1+\dfrac{0.06}{12}\right)^{-300}}{\dfrac{0.06}{12}} = 459784.81\) \(2962.40\cdot\dfrac{1-\left(1+\dfrac{0.06}{12}\right)^{-300}}{\dfrac{0.06}{12}} = 459121.34\) \(459784.81-459121.34 = 663.47\) \(2962.40 - 663.47 = 2298.93\) This suggests that I start this conversation over!! Have you considered calculating it directly? $460,000*(0.06/12) = $2,300.00 -- So much better! Sorry about the false start.

OpenStudy (anonymous):

Yes! I got 2,300 but then i didn't know what to do from there. $2,300 is the total interest correct? well I thought that if I took that and divided by 300 it would be right but it's not

OpenStudy (anonymous):

I got 300 by multiplying 25 and 12 since there is a 25 year period and there are 12 months.... @tkhunny

OpenStudy (anonymous):

(in a year)

OpenStudy (tkhunny):

All you need is this: $460,000*(0.06/12) = $2,300.00 - Calculate the interest directly.

OpenStudy (anonymous):

So you're saying that the answer would be 2,300? Cause I tried that too.. :l

OpenStudy (tkhunny):

The WAY LONGER way was only $1.07 different. Let's go with it calculated directly at $2,300.00.